If sec a.sin(36+a)=1 then find the value of a so that a and 36 +a are acute angles.Find a if sin 5a = cos 4a where 4a and 5a are acute angles. I will be highly obliged to for both the answers.

Grade:10

1 Answers

Utsav Basu
6 years ago
Ans.: 1)   
     sec a * sin (36 + a) = 1
or, sin (36 + a) = cos a
or, sin (36 + a) = sin (90 – a)
or, 36 + a = 90 – a
or, a = 27
 
Ans.: 2)
     sin 5a = cos 4a
or, sin 5a = sin (90 – 4a)
or, 5a = 90 – 4a
or, a = 10
 
I hope I am able to answer your questions satisfactorily. Please approve my answers if so...